### v1

# CraftRTL: High-quality Synthetic Data Generation for Verilog Code Models with Correct-by-Construction Non-Textual Representations and Targeted Code Repair

Abstract
Despite the significant progress made in code generation with large language models, challenges persist, especially with hardware description languages such as Verilog. This paper first presents an analysis of fine-tuned LLMs on Verilog coding, with synthetic data from prior methods. We identify two main issues: difficulties in handling non-textual representations (Karnaugh maps, state-transition diagrams and waveforms) and significant variability during training with models randomly making “minor” mistakes. To address these limitations, we enhance data curation by creating correct-by-construction data targeting non-textual representations. Additionally, we introduce an automated framework that generates error reports from various model checkpoints and injects these errors into open-source code to create targeted code repair data. Our fine-tuned Starcoder2-15B outperforms prior state-of-the-art results by 3.8%, 10.9%, 6.6% for pass@1 on VerilogEval-Machine, VerilogEval-Human, and RTLLM.
1 Introduction
Large Language Models (LLMs) have achieved significant success across various natural language processing tasks and have extended their capabilities to code generation, leading to the development of specialized models targeting code generation. The effectiveness of these models is largely influenced by the size and quality of their training datasets, as highlighted by scaling laws (Achiam et al., 2023; Zhang et al., 2024a). Prominent code LLMs have set new benchmarks records by utilizing extensive, synthetically generated datasets through methods like Self-Instruct (Wang et al., 2022; Chaudhary, 2023), Evol-Instruct (Xu et al., 2023), and OSS-Instruct (Wei et al., 2023). These synthetic data generation techniques allow code LLMs to generate a wide range of complex code examples, enhancing their training and performance in real-world coding scenarios.
While most code LLMs concentrate on software programming languages, there is increasing interest in developing models for hardware description languages (HDLs), which are essential for chip design and hardware verification. Despite efforts to collect and synthesize more diverse Verilog code to enhance specialized code LLMs (Liu et al., 2023c; Pei et al., 2024; Cui et al., 2024; Zhao et al., 2024), HDLs still face challenges akin to those encountered in low-resource languages (Cassano et al., 2022). These challenges are mainly due to the limited availability of high-quality instruction-following data and the constrained capability of existing LLMs to generate RTL code, which affects the models’ performance and their ability to generalize across programming languages.
Developing high-quality synthetic Verilog code for training code large language models (LLMs) faces significant challenges due to two primary factors. Firstly, Verilog is considered a low-resource language (Cassano et al., 2022), meaning there is a scarcity of available training data compared to high-resource software programming languages like Python. This limited data availability restricts the models’ ability to learn diverse and complex coding patterns effectively. Secondly, verifying the correctness of hardware description language (HDL) code, such as Verilog, is inherently more complex than verifying software code. While software code correctness can often be assessed using random test cases and automated unit tests (Chen et al., 2022), hardware code requires comprehensive testbenches and rigorous verification planning and methodologies. This additional complexity makes it challenging to ensure that synthetic Verilog code is functionally accurate (Bhandari et al., 2024; Qiu et al., 2024), posing a barrier to improving model performance.
In this paper, we start with a thorough analysis of fine-tuned large language models (LLMs) applied to Verilog code, using synthetic data techniques from previous works. Our analysis reveals two key issues: (1) models have difficulty handling non-textual elements in problem statements, indicating challenges in interpreting complex or unconventional inputs; and (2) there is notable variability in the models’ pass rates across different benchmark problems and training checkpoints, exposing inconsistencies in learning outcomes, often due to the models making “minor” programming mistakes.
Given the limitations identified in our analysis of relying solely on LLMs for generating synthetic data, we shift our focus to improving data curation to address these issues. Current LLMs frequently struggle with interpreting and processing non-textual representations and are insufficient in generating effective testbenches for evaluating solution quality. Therefore, instead of depending exclusively on LLMs to address data quality concerns, we develop targeted fine-tuning data to better mitigate these problems. Experimental results demonstrate that our models achieve state-of-the-art (SOTA) results on VerilogEval (Liu et al., 2023b) and RTLLM v1.1 (Lu et al., 2024) benchmarks, outperforming prior works by large margins on problems with human-level description. The major contributions of this paper are as follows:
  • We perform a thorough analysis of fine-tuned LLMs on Verilog code using previously established synthetic data generation methods, uncovering challenges with non-textual elements and notable variability in performance across benchmark problems during training.
  • We create correct-by-construction data to ensure solution correctness, incorporating Karnaugh Maps, state-transition diagrams, and waveforms, which significantly enhance the model’s ability to handle non-textual representations.
  • We develop an automated framework that utilizes LLMs to generate error reports from benchmark problems at various checkpoints, which are then injected into open-source code to create a fine-tuning dataset targeted at correcting the model’s specific “minor” mistakes.
  • We rigorously evaluate the latest foundational and frontier code models. We note that recent advanced models like GPT-4o already reached competitive performance compared to previous efforts targeting Verilog code generation.
  • Experimental results demonstrate that models fine-tuned with our data achieve state-of-the-art performance on Verilog coding. Specifically, our fine-tuned model based on Starcoder2-15B (Lozhkov et al., 2024) outperforms prior SOTA results by 3.8%, 10.9%, 6.6% for pass@1 on VerilogEval-Machine, VerilogEval-Human, and RTLLM, respectively.
2 Examining fine-tuned LLMs Using Synthetic Generated Data on Verilog Coding
In this section, we start with a thorough analysis of fine-tuned large language models (LLMs) applied to Verilog code. We adapt previous approaches for generating synthetic data for general coding to focus on Verilog code. For our pilot study, we only present results based on fine-tuning StarCoder2-15B (Lozhkov et al., 2024). Details on experimental settings are the same as in Section 4. We assess model performance in Verilog code completion and identify two main issues. First, the models demonstrate notably poor performance when dealing with non-textual elements in the problem statements. Second, the variability in the models’ pass rates across different benchmark problems and training checkpoints suggests inconsistencies in learning outcomes and model variability.
2.1 Synthetic Data Generation for Verilog Coding
We build on previous methods for synthetic data generation by applying Self-Instruct (Wang et al., 2022) and OSS-Instruct (Wei et al., 2023) with custom prompt templates tailored for Verilog coding. To enhance data coverage and diversity, we supplement these techniques with additional context from Wikipedia and textbooks. We also prompt models to generate problem descriptions to include non-textual representations.
Table 1: Data quantity SDG.
Method Quantity
Self-Instruct 24.7k
OSS-Instruct 28.4k
Docu-Instruct 12.0k
Non-textual 15.0k
SDG Total 80.1k
We use nemotron-4-340b-instruct (Nvidia et al., 2024) selected for its open license that allows commercial use. Our process includes deduplication and a decontamination procedure akin to that outlined by Li et al. (2023). Additionally, we conduct syntax checks to eliminate coding problems containing docstrings or solutions from Verilog benchmarks. To ensure further data quality, we discard code solutions that fail these syntax checks and apply self-verification (Weng et al., 2023) to remove entries where the LLM identifies errors in the solution. Table 1 shows the quantity of our synthetic data generation (denoted as SDG) after deduplication and filtering, yielding a total of 80.1k fine-tuning examples.
Self-Instruct
We follow the approach outlined in Wang et al. (2022) to generate synthetic Verilog coding problems. Initially, we randomly generate from the LLM and curate 50 questions that request Verilog coding problems without any in-context examples. From these, we then randomly choose 1 to 5 seed questions to use as in-context examples.
OSS-Instruct
We begin by processing pretraining code data to extract our seed code from The Stack v2 (Lozhkov et al., 2024), focusing on Verilog and SystemVerilog. Following the approach in Liu et al. (2023b), we post-process this data by selecting self-contained Verilog code that passes syntax checks using Pyverilog (Takamaeda-Yamazaki, 2015). With the refined seed code data, we then prompt large language models (LLMs) to use this code as inspiration for generating Verilog coding problems similar to Wei et al. (2023).
Docu-Instruct
Drawing inspiration from Nvidia et al. (2024) and Sudalairaj et al. (2024), we utilize document sources from Wikipedia and textbooks for instruction generation. We begin by filtering Wikipedia entries, prompting the LLM to classify whether the content pertains to hardware design or Verilog coding concepts. Additionally, we manually selected approximately relevant 100 textbooks. These textbooks are then segmented into chunks of paragraphs or sentences, ensuring each chunk contains fewer than 2k tokens.
Non-textual Representations
VerilogEval-Human (Liu et al., 2023b) includes benchmark problems involving non-textual representations. For example, Boolean logic tables and Karnaugh maps are presented in tabular formats, state-transition diagrams for finite state machines are depicted as edge lists and sequential waveforms are described in tables with signal outputs recorded at various time steps. To incorporate such representations, we encouraged LLMs to generate problems from open-source code, with instructions to utilize these tabular data structures.
2.2 Challenges with Non-Textual Representations
Table 2: pass@1 results on VerilogEval sampled with temperature of 0.8.
Model Machine Human NonText
GPT-4o 63.7 55.4 27.0
Starcoder2 57.7 29.1 10.3
Starcoder2-SDG 73.7 47.4 22.2
We observe that models underperform on benchmark problems involving non-textual input formats, such as Karnaugh Maps, state-transition diagrams, and waveforms. Table 2 shows the pass@1 results for the VerilogEval (Liu et al., 2023b). Additionally, we have identified a subset of 45 questions within VerilogEval-Human that include non-textual representations, termed VerilogEval-NonText. It appears that models like GPT-4o and Starcoder2 struggle with these non-textual formats, likely due to insufficient representation of such data during both pretraining and fine-tuning. Despite our efforts to generate such questions during synthetic data creation, our fine-tuned models still lag in these areas. This outcome is not entirely surprising, given that the LLMs used were also ineffective at generating problems with these representations, complicating the validation of fine-tuning data. These results suggest that merely including non-textual data is insufficient; ensuring the quality and correctness of the data, particularly that the code solutions accurately align with these representations, is crucial.
2.3 Variability on Pass Rates During Training
Figure 1: Agreement by pass rate between two subsequent checkpoints during training on SDG data for problems in VerilogEval-Human.
During our training, we observed significant variability in the model’s pass rate on specific benchmark problems across different checkpoints. We note such variance is different from training instability (Wortsman et al., 2023) as we observe a stable decrease in the training loss. This variability persists even in the later stages of training, despite using a low learning rate. We illustrate this variability in Figure 1. The scatter plot tracks the pass rate for each problem in VerilogEval-Human, with each point representing the pass rate for the same problem across two checkpoints. The size of each point indicates the number of problems with the same pass rates for the two model checkpoints. We further categorize the region into areas where the checkpoints agree on problem difficulty and areas where they do not.
Alarmingly, we find that nearly 15% of the problems show significant discrepancies between these two checkpoints, with an equal number of problems demonstrating improvement and degradation. Our detailed analysis of the sampled code completions for such problems when pass rate degrades suggests that the model is generally on the right track but makes “minor” errors that are small, detailed, and seemingly trivial. While it is possible that LLMs experience catastrophic forgetting during fine-tuning (Luo et al., 2024a), we do not anticipate this being a major factor due to the low learning rate and the small number of gradient updates (64 steps with 16k data samples). Instead, we believe the primary issue is our inability to ensure the quality of our data, particularly in verifying whether the sampled code solutions correctly solve the code problems.
3 Improving Verilog Coding with Correct-by-Construction Non-Textual Representations and Targeted Code Repair
Based on our detailed analysis of the limitations of relying solely on LLMs for generating synthetic data, we focus our data curation efforts to address these shortcomings. Our goal is to enhance data quality and ensure the correctness of solutions for the generated problems. We have found that current LLMs often lack the capability to understand and process non-textual representations effectively and are unable to generate satisfactory testbenches for assessing solution quality. Consequently, rather than depending entirely on LLMs to resolve data quality issues, we instead create targeted fine-tuning data to mitigate these problems.
Figure 2: Generating non-textual problems and correct-by-construction solutions.
3.1 Ensuring Quality Through Correct-by-Construction
Table 3: Data quantity CC.
Method Quantity
KMap 12.5k
FSM 8.0k
Waveforms 8.0k
CC Total 28.5k
We generate Verilog code problems and solutions that are correct-by-construction. Our focus is on creating problems and solutions for non-textual representations. The overview of the process for generating non-textual problems and correct-by-construction solutions is illustrated in Figure 2. Table 3 shows the quantity of our correct-by-construction generation data (referred to as CC). To prevent data contamination, we exclude entries that duplicate the data representations of benchmark problems.
Karnaugh Maps and Truth Tables (KMap)
We start by sampling random configurations, which include selecting the number of variables and their names. After determining the number of variables, we randomly choose valid minterms and don’t-cares. For $n$ variables, there are $2^{n}$ possible states, and each state can be assigned one of three values (0, 1, or x), leading to $3^{2^{n}}$ possible combinations of minterms and don’t-cares. From these minterms, we derive the sum-of-products (SOP) form to represent the Boolean logic. We then create Truth Tables and Karnaugh Maps based on the chosen minterms and don’t-cares. In the KMap, Gray encoding is used as default for the row and column sequences to ensure that only a single bit changes between adjacent cells. Additionally, we apply modifications by transposing the map and randomly swapping adjacent rows or columns. We randomly sample from $n=\{3,4\}$ variables.
State Transition Graphs and Tables (FSM)
We construct problems for finite-state machines (FSMs) with state-transition representations with a similar approach to KMaps. We begin by sampling random configurations, including the number of states (e.g., 4, 6, or 10) and the bit width of the input (e.g., 1 or 2). We then create the transition graph, ensuring that it is both meaningful and legally defined. We generate state-transition graphs for both Moore and Mealy state machines. From these graphs, we produce edge-list and transition table representations. Finally, we construct the Verilog code to implement the logic for state transitions and output assignments.
Algorithm 1 Generate transition graph for Moore FSM.
  Input: Number of states $n$, bit width of input $w$
  Output: FSM graph with transitions and states
  Initialize the number of states $n$ and bit width of input $w$
  Randomly generate a tree with $n$ nodes
  Define the root of the tree as the reset state
  for each node in the tree do
     Assign a unique state to the node
     Assign an output to the node
  end for
  for each node in the tree do
     Add additional transition edges to form a graph
     Ensure that each node has an out-degree of $2^{w}$
  end for
Algorithm 1 outlines the process for generating a Moore FSM with random transitions. State reachability is ensured by first constructing a tree. Legality for state transition is ensured by ensuring each node has an out-degree of $2^{w}$ with the input bit width of $w$. The result is an FSM where the transitions between states are randomly assigned but conform to the specified input bit width. The algorithm can be easily modified for a Mealy FSM by assigning the output to the edges rather than nodes.
Figure 3: State transition logic.
Figure 3 illustrates our approach for generating state transition logic in Verilog from a state-transition graph. Our method predominantly employs an out-edge focused strategy for state transitions. Additionally, we incorporate in-edge focused transition logic to address specific challenges encountered in benchmark problems. These benchmarks often involve states represented using one-hot encoding and require rigorous testing of non-default states.
Waveforms
We utilize correct-by-construction code solutions for both KMaps and FSMs. Because these codes are generated using similar templates, designing corresponding testbenches is straightforward. We simulate the generated code to produce waveform Value Change Dump (VCD) files. These VCD files are then parsed and converted into waveform representations. Our approach covers KMaps as combinational circuits and FSMs as sequential circuit waveforms.
3.2 Mitigating “Minor” Errors with targeted code repair
Figure 4: An overview of our approach in generating targeted code repair data.
Our analysis revealed that the models were generally on the right track to correct solutions but were making minor errors—small, detailed, and seemingly trivial. Unlike complex, unsolvable problems, these minor errors could be easily corrected by language models. This insight led us to develop a new strategy centered on targeted error code repair. Our approach includes creating detailed error reports on benchmark problems, re-creating these errors on correct open-source code, and conducting rigorous validation to ensure quality. We use nemotron-4-340b-instruct as the LLM to construct our targeted Repair data. We generated 847 error reports across the three benchmarks and produced 2,736 data samples. After filtering, this resulted in a final set of 1,406 targeted code repair data points.
Error Report Construction
To systematically address the issue, we first created a comprehensive Error Report for benchmark problems using LLMs, targeting those with significant pass rate fluctuations across training checkpoints for models on SDG data. We prompt the LLM to examine the nature of the mistakes by comparing correct and erroneous code completions for each problem, categorizing the errors into functional types (such as logic, multiplexing, and arithmetic operations). Additionally, we prompt LLMs to identify mistakes in common error patterns, for example, frequent mistakes in arithmetic shifts, bit manipulation, and unintended latch creation. This detailed report not only categorizes the errors but also highlights areas where the model consistently underperforms.
Targeted Code Repair Dataset
Building on the error report, we further develop a targeted code repair dataset to address these common errors. This dataset is constructed using two main sources: the errors identified in the Error Report and correct code snippets gathered from open-source repositories. We introduced the identified errors into correct code snippets to create repair problems, which include a problem description, erroneous code implementation, and hints about the nature of the error and how to fix it. This targeted strategy enables the model to learn how to avoid common errors and generate improved code completions, thereby enhancing model accuracy.
Quality Assurance with LLM Validation
To ensure the reliability of the error report and the code repair dataset, we implemented a two-phase validation process with LLMs. In the first phase, we conducted a self-consistency check of the Error Report by having the language model attempt to fix the error code based on the report’s hints. This step verifies the accuracy of the report by confirming that the model can resolve the errors using the provided guidance. In the second phase, during the generation of the code repair dataset, we apply self-verification, including deduplication, syntax filtering, and benchmark decontamination. These measures ensure the dataset’s quality and uniqueness, preventing overlap with evaluation benchmarks.
4 Experiment
4.1 Implementation Details
Training Data
Our fine-tuning training data is comprised of 80.1k LLM synthetic generated data using various prompting methods as described in Section 2.1, 28.5k data samples generated correct-by-construction aimed at non-textual representations detailed in Section 3.1, and 1.4k carefully filtered data for targeted code repair as outlined in Section 3.2. We refer to each data set as SDG, CC, and Repair, respectively.
Pretrained Models
Following prior work, we use CodeLlama-7b-Instruct (Roziere et al., 2023) and Deepseek-Coder-6.7b-Instruct (Guo et al., 2024) as the base model, formatting our data according to their default chat prompt templates. Additionally, we explore the Starcoder2-15B (Lozhkov et al., 2024) model in our experiments.
Model Training
Training is conducted with 32 NVIDIA A100-80GB GPUs through the Distributed Data Parallel (DDP) module from PyTorch. We set the learning rate at 5e-5 for CodeLlama and DeepSeek-Coder, and 1e-5 for Starcoder2. We use Adam (Kingma & Ba, 2017) as our optimizer with full parameter updates and truncate sequence lengths longer than 4096 tokens. We used a batch size of 256 samples. We fine-tune models for 1 epoch using a standard cross entropy loss on the response tokens (while masking loss on prompt tokens).
Model Inference
We use vLLM (Kwon et al., 2023) where the inference engine is set up with bf16 dtype, tensor parallel size of 8, and a maximum token limit of 4096. We sample each problem 20 times. We report the best results from two different temperatures 0.2 and 0.8, as consistent with prior work (Liu et al., 2023c; Zhao et al., 2024).
Table 4: We compare our models with various baseline models on VerilogEval (Liu et al., 2023b). We update the results from Zhao et al. (2024) with the latest foundational and frontier code models. The best results for each model type are highlighted in bold. We use red to denote the best model overall and blue to indicate the best results among models other than ours.
Type Model Size VerilogEval (Liu et al., 2023b)
Machine (%) Human (%)
pass@1 pass@5 pass@10 pass@1 pass@5 pass@10
Foundational Models Llama-3.1 8B 48.7 67.3 74.1 26.9 37.8 44.2
Llama-3.1 405B 67.3 75.1 76.9 53.8 61.0 62.8
Nemotron-4 340B 53.0 60.3 62.2 43.1 48.3 50.0
GPT-3.5-turbo - 58.0 74.0 77.6 31.2 44.1 47.4
GPT-4o - 65.9 71.4 72.7 57.1 63.9 66.7
Code Models CodeLlama 7B 43.1 47.1 47.7 18.2 22.7 24.3
CodeQwen 7B 46.5 54.9 56.4 22.5 26.1 28.0
Starcoder2 15B 68.7 82.3 88.5 37.7 50.6 57.2
DeepSeek-Coder 6.7B 52.2 55.4 56.8 30.2 33.9 34.9
DeepSeek-Coder-V2 16B 67.4 78.3 81.8 46.9 55.9 58.9
DeepSeek-Coder-V2 236B 68.2 74.1 76.2 56.4 62.2 66.0
RTLCoder (Liu et al., 2023c) Mistral 7B 62.5 72.2 76.6 36.7 45.5 49.2
DeepSeek-Coder 7B 61.2 76.5 81.8 41.6 50.1 53.4
BetterV (Pei et al., 2024) CodeLlama 7B 64.2 75.4 79.1 40.9 50.0 53.3
DeepSeek-Coder 6.7B 67.8 79.1 84.0 45.9 53.3 57.6
CodeQwen 7B 68.1 79.4 84.5 46.1 53.7 58.2
CodeV (Zhao et al., 2024) CodeLlama 7B 78.1 86.0 88.5 45.2 59.5 63.8
DeepSeek-Coder 6.7B 77.9 88.6 90.7 52.7 62.5 67.3
CodeQwen 7B 77.6 88.2 90.7 53.2 65.1 68.5
OriGen (Cui et al., 2024) DeepSeek-Coder 6.7B 74.1 82.4 85.7 54.4 60.1 64.2
Ours SDG-CC-Repair CodeLlama 7B 78.1 85.5 87.8 63.1 67.8 69.7
DeepSeek-Coder 6.7B 77.8 85.5 88.1 65.4 70.0 72.1
Starcoder2 15B 81.9 86.9 88.1 68.0 72.4 74.6
4.2 Evaluation Metric and Benchmark
Evaluation Metric Following prior work (Chen et al., 2021; Liu et al., 2023a), for each experiment we use the unbiased pass@k metric to measure the Verilog generation accuracy. The pass@k metric estimates the proportion of problems that can be solved at least once in k attempts:
$pass@k:=\mathbb{E}_{\text{Problems}}\left[1-\frac{\binom{n-c}{k}}{\binom{n}{k}}\right],$ (1)
where $n\geq k$ represents the total number of trials for each problem, and $c$ represents the number of trials that pass the functional check.
VerilogEval (Liu et al., 2023b) contains two subsets of problems, where VerilogEval-Human contains manually converted problem descriptions from the original HDLBits website, and VerilogEval-Machine with GPT-3.5 generated problem descriptions.
RTLLM (Lu et al., 2024) is an open-source benchmark designed for generating Register Transfer Level (RTL) code from natural language instructions. It evaluates models on syntax correctness, functional correctness, and design quality, offering a thorough analysis of model outputs.
4.3 Results
Table 5: Evaluations on RTLLM v1.1 (Lu et al., 2024) using unbiased pass@k metrics. The best results for each model type are highlighted in bold. We use red to denote the best model overall and blue to indicate the best results among models other than ours. We re-evaluate all models (see Appendix A for details).
Type Model Size RTLLM v1.1 (Lu et al., 2024)
Syntax (%) Func. (%)
pass@1 pass@5 pass@10 pass@1 pass@5 pass@10
Foundational Models Llama-3.1 8B 40.7 60.6 65.5 19.3 34.7 37.9
Llama-3.1 405B 56.5 64.4 72.4 38.9 45.8 51.7
Nemotron-4 340B 41.7 47.2 48.3 18.9 20.7 20.7
GPT-3.5-turbo - 50.3 61.2 65.5 28.3 36.9 41.4
GPT-4o - 50.3 59.9 62.1 33.8 44.4 48.3
Code Models CodeLlama 7B 46.6 62.6 68.9 17.9 29.9 34.5
CodeQwen 7B 45.8 65.8 72.4 24.1 34.0 37.9
Starcoder2 15B 38.3 81.0 94.7 15.5 37.6 45.7
DeepSeek-Coder 6.7B 51.4 64.4 68.9 23.1 29.3 34.5
DeepSeek-Coder-V2 16B 51.4 57.8 58.6 33.1 37.1 37.9
DeepSeek-Coder-V2 236B 63.4 78.1 79.3 34.5 50.2 55.1
RTLCoder (Liu et al., 2023c) Mistral 7B 64.6 73.7 78.3 24.5 37.3 42.3
DeepSeek-Coder 6.7B 73.4 83.9 86.2 35.8 40.3 43.1
CodeV (Zhao et al., 2024) CodeLlama 7B 79.0 89.2 89.9 39.4 50.3 53.1
DeepSeek-Coder 6.7B 78.3 87.4 89.1 42.4 51.5 53.2
CodeQwen 7B 78.8 89.5 92.4 36.6 53.3 61.3
OriGen (Cui et al., 2024) DeepSeek-Coder 6.7B - - - - 65.5 -
Ours SDG-CC-Repair CodeLlama 7B 85.7 93.9 94.8 42.6 52.9 58.2
DeepSeek-Coder 6.7B 84.3 92.9 95.4 53.1 58.8 62.6
Starcoder2 15B 79.8 93.9 96.2 49.0 65.8 74.5
Main Results
Table 4 and Table 5 compare our models with baselines on VerilogEval and RTLLM. We mainly source baseline results from Zhao et al. (2024). For RTLLM we found a large variance with biased pass@5, thus we re-evalaute all models and report unbiased pass@k metric. We further rigorously evaluate the latest foundational and frontier code models, including Llama-3.1 (Dubey et al., 2024), DeepSeek-Coder-V2 (DeepSeek-AI et al., 2024), and GPT-4o. Recent foundational and frontier code models already reached competitive performance compared to previous efforts targeting Verilog code generation.
Compared to previous approaches like CodeV (Zhao et al., 2024), our models achieve comparable performance on VerilogEval-Machine and show significant improvements on benchmarks with human-like descriptions. Machine descriptions often provide detailed, line-by-line coding instructions, whereas human descriptions are high-level, integrating problem-solving skills and a deeper understanding of the hardware module’s functionality. Enhancing the model’s ability to handle human-like descriptions is crucial, as these more accurately reflect how designers interact with the models and set expectations for Verilog generation. Our fine-tuned Starcoder2-15B surpasses previous state-of-the-art results by 3.8%, 10.9%, and 6.6% in pass@1 metrics on VerilogEval-Machine, VerilogEval-Human, and RTLLM, respectively.
Table 6: Ablation study on training data.
Model VerilogEval RTLLM v1.1
Machine Human Func
pass@1 (%) pass@5 (%)
Starcoder2-15B 68.7 37.7 37.6
SDG 75.2 54.7 62.1
SDG-CC 73.9 62.0 62.8
SDG-CC-Repair 81.9 68.0 65.8
Table 6 highlights the effectiveness of our generated data fine-tuned on Starcoder2-15B. Our CC data enhances the model’s ability to handle non-textual representations, leading to improved scores on VerilogEval-Human. Our targeted code Repair data boosts performance across all benchmarks, suggesting that the model has learned to generalize from code repair tasks and reduce similar errors during code completion.
Improved Variability During Training
Figure 5: Agreement by pass rate between two subsequent checkpoints during training on SDG-CC-Repair for problems in VerilogEval-Human.
Figure 5 displays the pass rates for two consecutive checkpoints of Starcoder2-SDG-CC-Repair on VerilogEval-Human problems, sampled with a temperature of 0.8. Compared to Figure 1, the updated model shows significant improvements by (1) moving previously unsolved problems into the solved category, including those with non-textual representations addressed by our correct-by-construction CC data, and (2) reducing the number of problems with large pass rate discrepancies, particularly where performance had degraded. The targeted repair data has effectively mitigated the model’s tendency to repeat common mistakes found in our Repair dataset, despite the noise inherent in synthetically generated SDG data.
Scaling Data for Non-textual Representations
Figure 6: pass@1 on non-textual problems with total number of CC data with temperature 0.8.
Figure 6 illustrates the scaling of correct-by-construction (CC) data and the fine-tuned Starcoder2-15B pass rate on problems involving non-textual representations. We expanded our testing to include strictly in-distribution test set, with each category containing around 50 problems. The results show that the model can quickly learn and comprehend these non-textual representations with as few as 4k training data samples, with the pass rate steadily improving as more data is provided. Additionally, the model demonstrates the ability to generalize to VerilogEval-NonText benchmark problems. While our models achieve near-perfect scores on KMap and FSM problems, they perform less effectively on Waveforms, suggesting that reverse engineering circuits from waveforms pose a greater challenge.
Ensuring Quality for Targeted Code Repair
Table 7: Ablation study on Repair data quality with Starcoder2-15B.
Model VerilogEval RTLLM v1.1
Machine Human Func
pass@1 (%) pass@5 (%)
SDG-CC 73.9 62.0 62.8
SDG-CC-Repair 81.9 68.0 65.8
w/o self-consistency 75.3 63.3 63.7
w/o error report 76.9 59.6 59.4
The quality control mechanisms integrated into the data generation pipeline are crucial for improving model performance, particularly in correcting minor errors through targeted code repair. To evaluate the impact of these quality controls, we conducted an ablation study in Table 7), where we systematically removed each component of the targeted code repair generation pipeline and assessed the resulting model performance. Specifically, we eliminated the self-consistency checks that validate whether the generated error report effectively guides the LLMs in correcting mistakes. Additionally, we tested the removal of the error report entirely, substituting it with random errors injected into the open-source code by the LLMs. The benchmark results indicate a significant performance drop when these validation processes are excluded. These findings highlight the essential role of both the self-consistency checks and the targeted error report in improving the model’s ability to correct errors.
5 Related Work
Synthetic Data Generation for Model Fine-tuning.
The performance of large language models (LLMs) hinge on the quality and diversity of their training data. To address the limitations of manual datasets, synthetic data generation methods (Wang et al., 2022; Xu et al., 2023) have been developed to automatically create instruction-following examples from LLMs, reducing reliance on human annotations. Various techniques enhance data quality:  Wang et al. (2022) generates multiple reasoning traces and selects the most frequent output to improve robustness, while other approaches (Lightman et al., 2023; Zhang et al., 2024b) assess response quality based on these traces. Self-training methods utilize synthetic data for iterative fine-tuning, boosting reasoning capabilities (Singh et al., 2023; Feng et al., 2023). These advancements show how synthetic data can effectively scale and optimize models through iterative feedback.
Large Language Models for Code Generation.
Recent breakthroughs in large language models (LLMs) have greatly enhanced their capability to tackle complex code generation tasks. Much of the research focuses on developing LLMs specialized for code by continuing their pretraining on code data (Guo et al., 2024; Bai et al., 2023; Roziere et al., 2023; DeepSeek-AI et al., 2024) from open-source repositories like GitHub (Kocetkov et al., 2022; Lozhkov et al., 2024) and commit histories (Muennighoff et al., 2023). Further improvements to these models come from reinforcement learning (Le et al., 2022) and more often instruction fine-tuning, which involves techniques to address more complex coding problems (Luo et al., 2024b), increasing diversity with unlabeled open-source code (Wei et al., 2023; Yu et al., 2024; Wu et al., 2024), ensuring solution correctness through self-written tests (Chen et al., 2022), and validating and debugging code execution through interactions with LLM agents (Lei et al., 2024).
Large Language Models for Verilog Coding.
While most code LLMs target software languages, there is increasing interest in models for hardware description languages like Verilog, essential for chip design and verification (Liu et al., 2024). Previous work has addressed the challenge of limited data through various methods, including synthetic data generation (Liu et al., 2023c), multi-level summarization of open-source Verilog code (Zhao et al., 2024), and enhanced code augmentation with self-reflection based on compiler feedback (Tsai et al., 2023; Cui et al., 2024). Other approaches focus on improving functional correctness and circuit performance through Monte Carlo Tree Search (DeLorenzo et al., 2024) and discriminator-guided sampling (Pei et al., 2024).
6 Discussions
In this work, we refer to synthetic data generation as methods of using large language models (LLMs) in data generation. While our approach—ensuring correctness through correct-by-construction—could also be considered “synthetic” and resembles methods explored in works like AlphaGeometry (Trinh et al., 2024), our problems are much simpler and on a smaller scale. Our observations about the variability of models on specific problems align with the findings of  Meta AI (2024), where “the model knows how to produce the right answer, but it does not know how to select it.” Instead of striving for absolute data correctness, preference learning (Rafailov et al., 2024; Ethayarajh et al., 2024) or reinforcement learning (Bai et al., 2022; Le et al., 2022), we generate targeted repair data by analyzing errors and re-create such scenarios by injecting similar errors into open-source code, somewhat analogous to how humans consolidate memories during sleep by integrating new information with past experiences (Walker & Stickgold, 2004; Stickgold, 2005).
7 Conclusion
This paper addresses key challenges in Verilog code generation with correct-by-construction data generation and targeted code repair data strategies. We identified significant issues with synthetic data generation, including difficulties with non-textual representations and variability in performance during training across benchmarks. To address these challenges, we generated data that is correct-by-construction and create targeted repair data by injecting errors to open-source code. Our approach led to substantial improvements, with models fine-tuned using our methods achieving state-of-the-art results on VerilogEval and RTLLM benchmarks. These advancements highlight the effectiveness of our strategies in enhancing model performance in Verilog code generation.
Appendix A Detailed Results
A.1 Our Models
We present our models’ results on Verilog benchmarks tested with temperatures 0.2 and 0.8. We ablate across different data blends, with SDG indicating using LLM synthetic generated data in Section 2.1, CC indicating correct-by-construction data targeting non-textual representations in Section 3.1, and Repair representing our targed code repair dataset in Section 3.2.
Our results for RTLLM use the open-source Icarus Verilog simulator [^1] https://github.com/steveicarus/iverilog to check syntax and functional pass rates. This might lead to lower pass rate scores compared to previous work that used Synopsys VCS, as Icarus Verilog does not support all syntax.
Table 8: Results for our models, across different dataset and temperature on VerilogEval.
Model Dataset Temperature VerilogEval (Liu et al., 2023b)
Machine (%) Human (%)
pass@1 pass@5 pass@10 pass@1 pass@5 pass@10
Starcoder2-15b SDG 0.2 75.2 79.2 80.1 54.7 60.1 61.2
0.8 73.7 84.0 86.1 47.4 61.9 64.8
SDG-CC 0.2 73.9 78.1 79.5 62.0 65.6 67.0
0.8 72.9 84.1 87.1 58.5 70.3 73.7
SDG-CC-Repair 0.2 81.9 84.2 85.0 68.0 71.7 72.0
0.8 78.1 86.9 88.1 64.1 72.4 74.6
DeepSeek-6.7b-Instruct SDG 0.2 73.4 77.8 78.9 48.3 53.2 54.5
0.8 71.4 82.5 85.4 44.0 58.1 62.3
SDG-CC 0.2 72.6 78.2 79.3 58.5 62.6 63.5
0.8 70.2 83.1 85.4 56.3 67.0 70.7
SDG-CC-Repair 0.2 77.8 82.7 83.4 65.4 67.7 68.2
0.8 75.2 85.5 88.1 61.6 70.0 72.1
CodeLlama-7b-Instruct SDG 0.2 74.5 77.9 78.8 45.3 50.3 51.5
0.8 71.2 82.6 85.1 42.6 55.6 59.0
SDG-CC 0.2 74.2 77.4 78.1 55.1 61.0 62.4
0.8 70.0 81.2 83.7 51.6 64.4 67.7
SDG-CC-Repair 0.2 78.1 81.5 81.7 63.1 66.2 66.8
0.8 73.7 85.5 87.8 58.1 67.8 69.7
Table 9: Results for our models, across different dataset and temperature on RTLLM.
Model Dataset Temperature RTLLM v1.1 Lu et al. (2024)
Syntax (%) Func. (%)
pass@1 pass@5 pass@10 pass@1 pass@5 pass@10
Starcoder2-15b SDG 0.2 78.1 86.5 90.1 49.0 60.4 66.3
0.8 77.1 89.0 94.1 43.8 62.1 68.0
SDG-CC 0.2 78.3 89.3 92.7 45.5 58.3 62.0
0.8 76.9 92.6 95.5 38.4 62.8 70.4
SDG-CC-Repair 0.2 79.8 87.9 90.5 49.0 59.1 62.6
0.8 79.3 93.9 96.2 45.3 65.8 74.5
DeepSeek-6.7b-Instruct SDG 0.2 79.3 86.8 90.5 40.3 45.9 49.6
0.8 76.6 92.5 96.2 40.0 53.8 63.6
SDG-CC 0.2 73.6 84.5 86.0 44.3 52.2 54.3
0.8 76.7 90.5 93.8 39.5 56.4 63.1
SDG-CC-Repair 0.2 84.3 92.2 93.0 53.1 58.8 60.3
0.8 80.0 92.9 95.4 45.5 57.9 62.6
CodeLlama-7b-Instruct SDG 0.2 74.0 82.5 86.8 30.0 33.9 35.8
0.8 70.9 89.1 94.5 34.0 47.2 52.8
SDG-CC 0.2 75.0 90.2 94.6 39.7 44.4 47.2
0.8 76.4 93.9 96.3 35.5 47.6 52.7
SDG-CC-Repair 0.2 85.7 93.9 94.8 42.6 49.4 51.2
0.8 80.3 93.9 94.8 36.9 52.9 58.2
A.2 Foundational and Frontier Code Models
We present detailed results on recent foundational and frontier code models. We also re-evaluate all models on RTLLM using unbiased pass@k metric.
Table 10: Results on foundational and code models on VerilogEval.
Type Model Size Temp VerilogEval (Liu et al., 2023b)
Machine (%) Human (%)
pass@1 pass@5 pass@10 pass@1 pass@5 pass@10
Foundational Models Llama-3.1 8B 0.2 48.7 66.2 70.6 26.9 36.9 40.4
0.8 42.1 67.3 74.1 23.0 37.8 44.2
Llama-3.1 70B 0.2 66.7 73.8 76.9 48.7 53.6 55.1
0.8 64.5 77.7 80.4 48.0 57.0 60.9
Llama-3.1 405B 0.2 67.3 72.8 74.1 51.9 57.0 58.9
0.8 66.4 75.1 76.9 53.8 61.0 62.8
Nemotron-4 340B 0.2 53.0 59.1 61.5 43.1 43.9 44.9
0.8 50.8 60.3 62.2 40.8 48.3 50.0
GPT-3.5-turbo - 0.2 58.0 66.4 68.5 31.2 39.4 41.7
0.8 56.6 74.0 77.6 28.9 44.1 47.4
GPT-4 - 0.2 53.2 63.7 66.4 36.1 43.5 46.2
0.8 35.3 53.4 58.9 35.2 53.4 58.9
GPT-4-turbo - 0.2 57.8 66.7 70.6 54.1 61.2 62.8
0.8 56.9 69.5 73.4 53.6 63.6 66.7
GPT-4o - 0.2 65.9 68.9 69.2 57.1 61.3 62.2
0.8 62.9 71.4 72.7 55.4 63.9 66.7
Code Models Starcoder2 15B 0.2 68.7 76.7 78.6 37.7 48.3 51.1
0.8 57.7 82.3 88.5 29.1 50.6 57.2
DeepSeek-Coder-V2 16B 0.2 67.4 74.6 76.2 46.9 53.3 54.5
0.8 65.6 78.3 81.8 46.3 55.9 58.9
DeepSeek-Coder-V2 236B 0.2 68.2 72.7 75.0 56.4 60.7 64.3
0.8 66.5 74.1 76.2 54.8 62.2 66.0
Table 11: Results on foundational and code models on RTLLM.
Type Model Size Temp RTLLM v1.1 (Lu et al., 2024)
Syntax (%) Func. (%)
pass@1 pass@5 pass@10 pass@1 pass@5 pass@10
Foundational Models Llama-3.1 8B 0.2 39.7 53.1 55.2 19.3 25.8 27.6
0.8 40.7 60.6 65.5 17.6 34.7 37.9
Llama-3.1 70B 0.2 47.9 51.7 55.2 34.1 34.5 34.5
0.8 48.9 57.6 58.6 29.6 31.0 31.0
Llama-3.1 405B 0.2 56.5 63.9 65.5 38.9 45.0 48.3
0.8 52.1 64.4 72.4 35.8 45.8 51.7
Nemotron-4 340B 0.2 41.7 47.2 48.3 14.1 15.5 17.2
0.8 41.7 46.3 48.3 18.9 20.7 20.7
GPT-3.5-turbo - 0.2 50.3 58.2 58.6 28.3 36.9 41.4
0.8 48.2 61.2 65.5 24.1 36.9 41.4
GPT-4 - 0.2 49.3 65.9 68.9 30.0 44.4 48.3
0.8 42.8 61.2 65.5 25.9 40.0 44.8
GPT-4-turbo - 0.2 38.9 44.8 48.3 27.2 35.1 37.9
0.8 40.3 48.8 51.7 27.5 40.2 44.8
GPT-4o - 0.2 50.3 59.9 62.1 33.8 44.4 48.3
0.8 47.5 63.2 66.7 31.3 44.1 48.3
Code Models CodeLlama 7B 0.2 46.6 62.6 68.9 17.9 29.9 34.5
0.8 34.8 59.7 68.9 13.4 25.9 31.0
CodeQwen 7B 0.2 45.8 55.8 58.6 24.1 33.1 37.9
0.8 45.5 65.7 72.4 22.4 34.0 37.9
Starcoder2 15B 0.2 38.3 77.5 86.3 15.5 37.6 44.6
0.8 31.6 81.0 94.7 11.0 34.2 45.7
DeepSeek-Coder 6.7B 0.2 51.4 62.6 65.5 23.1 26.8 27.6
0.8 49.7 64.4 68.9 21.0 29.3 34.5
DeepSeek-Coder-V2 16B 0.2 51.4 51.7 51.7 33.1 34.5 34.5
0.8 51.4 57.8 58.6 30.0 37.1 37.9
DeepSeek-Coder-V2 236B 0.2 63.4 73.0 79.3 34.5 44.9 52.9
0.8 61.8 78.1 79.3 32.9 50.2 55.1
A.3 Details on Evaluations
We format the prompt input as follows for VerilogEval, where the detail_description is the problem description (Machine or Human) and prompt field is the problem module header. We include module headers to avoid confusion on the signals naming.
[…]
An example of mux2to1 in VerilogEval-Human:
Create a 2-1 multiplexer. When sel=0, choose a. When sel=1, choose b.
module top_module (
input a,
input b,
input sel,
output out
We use similar templates for RTLLM v1.1, where we extract the top module header from the reference solution and provide it as input. Below is an example of adder_8bit:
Please act as a professional verilog designer.
Implement a module of an 8-bit adder with multiple bit-level adders in combinational logic.
Module name:
adder_8bit
Input ports:
a[7:0]: 8-bit input operand A.
b[7:0]: 8-bit input operand B.
cin: Carry-in input.
Output ports:
sum[7:0]: 8-bit output representing the sum of A and B.
cout: Carry-out output.
Implementation:
The module utilizes a series of bit-level adders (full adders) to perform the addition operation.
Give me the complete code.
module adder_8bit(
input [7:0] a, b,
input cin,
output [7:0] sum,
output cout);
We use default chat templates and default system prompts for open-source models tested. For GPT models from OpenAI, we use the following system prompt:
Please act as a professional verilog designer.
We post-process model responses to extract code. We extract content enclosed by triple backticks and remove the language identifier (Verilog). We then extract code enclosed in module and endmodule keywords with response.find(’module’) and response.rfind(’endmodule’). If the extracted code does not include a module header, the reference solution’s module header will be prepended. The code is then tested with the provided testbenches with the Icarus Verilog (iverilog) simulator to evaluate for syntax and functional correctness. This might lead to lower pass rate scores for RTLLM compared to previous work that used Synopsys VCS, as Icarus Verilog does not support all syntax.
A.4 VerilogEval-NonText
We select the following 45 problems from VerilogEval-Human that consists of non-textual representations in their problem descriptions:
2012_q1g, 2012_q2b, 2012_q2fsm, 2013_q2afsm, 2014_q3bfsm, 2014_q3c, always_nolatches, circuit1, circuit10, circuit2, circuit3, circuit4, circuit5, circuit6, circuit7, circuit8, circuit9, ece241_2013_q7, ece241_2014_q3, ece241_2014_q5b, fsm1, fsm1s, fsm2, fsm2s, fsm3, fsm3comb, fsm3onehot, fsm3s, fsm_onehot, fsm_ps2data, kmap1, kmap2, kmap3, kmap4, m2014_q3, m2014_q6, m2014_q6b, m2014_q6c, mt2015_q4, mt2015_q4a, mt2015_q4b, review2015_fsmonehot, rule110, rule90, truthtable1
A.5 Template Problems for Correct-by-Construction Data
When generating correct-by-construction CC data, we select 11 problems from VerilogEval-NonText to use as representative templates for constructing our prompts. To prevent contamination, we ensure that benchmark problems are excluded from our data. While our prompts will resemble those of the selected problems, the non-textual representations and solutions will differ. Additionally, to prevent overfitting to specific prompt templates, we use LLMs to rewrite the problem instructions for 20% of our data. Furthermore, we create validation test problems that are strictly in-distribution, based on the chosen problems.
Karnaugh Maps and Truth Tables: kmap1, m2014_q3, truthtable1.
State Transition Graphs and Tables: 2012_q2b, 2014_q3c, ece241_2014_q5b, fsm3comb, fsm3onehot, fsm_onehot, m2014_q6b, m2014_q6c.
Waveforms: We do not base our data on any benchmark problems specifically.
Table 12: Scaling Repair data.
Model VerilogEval RTLLM v1.1
Machine Human Func
pass@1 (%) pass@5 (%)
SDG-CC 73.9 62.0 62.8
SDG-CC-Repair 1k 81.9 68.0 65.8
SDG-CC-Repair 7k 82.2 67.4 64.5
A.6 Scaling Repair Data
As shown in Table 12, a carefully filtered dataset of 1.4k samples achieves comparable performance to a 7.8k dataset. This suggests that merely increasing the dataset size by injecting the same types of errors does not contribute meaningfully to improving model performance.
Table 13: Iterative code repair.
Model VerilogEval RTLLM v1.1
Machine Human Func
pass@1 (%) pass@5 (%)
SDG-CC 73.9 62.0 62.8
SDG-CC-Repair Iter 1 81.9 68.0 65.8
SDG-CC-Repair Iter 2 81.3 68.1 65.6
A.7 Iterative Code Repair
We conduct a second iteration by generating 2.7k repair data for the model based on the Repair data from the first iteration. As shown in  Table 13, performance mostly saturates after this initial iteration. We suspect that the remaining issues are likely due to significant errors that are challenging to correct.
A.8 Diversity of Generated Code
We assess the diversity of the code generated by our models. We measure this diversity using BLEU score, Jaccard similarity, and abstract tree edit distance (TSED) in Song et al. (2024). The VerilogEval-Human problems are categorized into NonText and Text, as described in Section A.4. For each problem, we compute the average code diversity score across sampled codes for the same problem and report the mean score for all problems. For TSED, we use PyVerilog (Takamaeda-Yamazaki, 2015) to extract the abstract syntax tree, and codes that fail syntax checks are excluded from the analysis.
Table 14 presents the results on code diversity. We sample 20 solutions with temperature of 0.8 for each model. We observe that fine-tuned models generally show a decrease in code diversity for both Text and NonText problems. This reduction is expected, as BLEU and Jaccard metrics account for both correct and incorrect code solutions, and there are often multiple ways to implement a correct solution. When comparing our fine-tuned models with GPT-4o, code diversity is similar for Text problems, but our models exhibit poor diversity for NonText problems. This is anticipated, given that the CC training dataset for NonText problems is generated using correct-by-construction methods and follows similar templates for Verilog code. However, our models demonstrate comparable diversity to GPT-4o for Text problems, particularly in TSED metric.
Table 14: Diversity of generated code solutions on VerilogEval-Human sampled with temperature of 0.8. Lower scores indicate higher diversity.
Type Models Text NonText
Jaccard BLEU TSED Jaccard BLEU TSED
Pretrained Models CodeLlama 0.5330 0.3808 0.4255 0.4707 0.2507 0.3521
DeepSeek-Coder 0.6606 0.5454 0.5956 0.6548 0.3797 0.3847
Starcoder2 0.7724 0.5084 0.5520 0.7212 0.3607 0.4020
GPT-4o 0.6798 0.6633 0.6906 0.7390 0.6376 0.6137
Ours SDG-CC-Repair CodeLlama 0.6848 0.5992 0.6354 0.8583 0.7242 0.7158
DeepSeek-Coder 0.6828 0.6040 0.6319 0.8308 0.6866 0.6598
Starcoder2 0.7018 0.6381 0.6721 0.8799 0.7750 0.7740
Type Models VerilogEval-Human (Overall)
Jaccard BLEU TSED
Pretrained Models CodeLlama 0.5155 0.3441 0.4156
DeepSeek-Coder 0.6590 0.4987 0.5505
Starcoder2 0.7580 0.4667 0.5198
GPT-4o 0.6965 0.6561 0.6802
Ours SDG-CC-Repair CodeLlama 0.7333 0.6345 0.6515
DeepSeek-Coder 0.7246 0.6273 0.6379
Starcoder2 0.7512 0.6767 0.6942
Appendix B Examples of Targeted Code Repair Data
Problem description: vector3 from VerilogEval-Machine The top module contains 6 ports that are fed into a concatenation which outputs 4 ports. The first input "a" is fed into the first port in the concatenation, the second input "b" is fed into the second port of the concatenation, the third input "c" is fed into the third port of the concatenation, the fourth input "d" is fed into the fourth port of the concatenation, the fifth input "e" is fed into the fifth port of the concatenation, the sixth input "f" is fed into the sixth port of the concatenation. The last bit of the concatenation is 2’b11. This value is fed in to the last port of the concatenation. module top_module ( input [4:0] a, input [4:0] b, input [4:0] c, input [4:0] d, input [4:0] e, input [4:0] f, output [7:0] w, output [7:0] x, output [7:0] y, output [7:0] z );
assign {w, x, y, z} = {a, b, c, d, e, f, 2’b11};
endmodule
Listing 1: Correct Code
// Concatenate the inputs with the last bit as 2’b11
assign {w, x, y, z} = {2’b11, a, b, c, d, e, f};
endmodule
Listing 2: Error Code
LLM generated Error Report Error Type: Incorrect vector concatenation and splitting Category: Combinatorial: wiring Description: Steps to repair the erroneous implementation: 1. Concatenating the two 1 bits at the beginning of the concatenated vector. 2. Assign the output vectors from the concatenated vector in the correct order and bit ranges For example: assign {w, x, y, z} = {2’b11, a, b, c, d, e, f}; To correct the implementation, you should concatenate the last bit as ‘2’b11’ on the right, as shown in the correct implementation: assign {w, x, y, z} = {a, b, c, d, e, f, 2’b11};
LLM constructed Repair data from open-source code Input: #### 1. Problem Description You are given a Verilog module that demonstrates the use of bit slicing and concatenation in a loop. The goal is to correctly display specific portions of a 32-bit register using a for-loop and explicit slicing. However, there is a commonly made error in the concatenation and bit slicing order that causes incorrect results to be displayed. Your task is to identify and fix the error in the code.
#### 2. Erroneous Implementation module block; reg [31:0] data; int i; initial begin data = 32’hFACE_CAFE; // Loop to display 8-bit segments of the data register for (i = 0; i < 4; i++) begin // Incorrectly reversing the bit order in the concatenation \$display ("data[8*%0d +: 8] = 0x%0h", i, data[8*(3-i) +: 8]); end // Incorrect slicing order, placing higher bits in lower indexes \$display ("data[7:0] = 0x%0h", data[31:24]); \$display ("data[15:8] = 0x%0h", data[23:16]); \$display ("data[23:16] = 0x%0h", data[15:8]); \$display ("data[31:24] = 0x%0h", data[7:0]); end endmodule In this erroneous implementation: - The loop erroneously reverses the order of the 8-bit slices. - The explicit bit-slicing outside of the loop has the slices ordered incorrectly, assigning the MSB to the least significant part and vice versa. #### 3. Hints for Fixing 1. **Correct the loop:** The loop currently reverses the order of the 8-bit slices. Instead of accessing ‘data[8*(3-i) +: 8]’, you should access ‘data[8*i +: 8]’ to display the correct slices of the register. 2. **Fix the slicing outside of the loop:** Ensure that the LSB (‘data[7:0]’) corresponds to the lower bits of the ‘data’ register, and the MSB (‘data[31:24]’) corresponds to the higher bits. Correct the slicing order for accurate display. The corrected implementation should ensure that both the loop and explicit bit-slicing access the register correctly, without reversing the bit order. Output: module block; reg [31:0] data; int i; initial begin data = 32’hFACE_CAFE; for (i = 0; i < 4; i++) begin \$display ("data[8*%0d +: 8] = 0x%0h", i, data[8*i +: 8]); end \$display ("data[7:0] = 0x%0h", data[7:0]); \$display ("data[15:8] = 0x%0h", data[15:8]); \$display ("data[23:16] = 0x%0h", data[23:16]); \$display ("data[31:24] = 0x%0h", data[31:24]); end endmodule
Figure 7: An example demonstrating the process for targeted code repair. In this example, training checkpoints during training would have the model sometimes generated correct and error solutions. We use LLMs to first summarize the errors for a detailed Error Report and then inject the errors to open-source code to construct Repair data.
Problem description: shift4 from VerilogEval-Human Build a 4-bit shift register (right shift), with asynchronous positive edge triggered areset, synchronous active high signals load, and enable. // (1) areset: Resets shift register to zero. // (2) load: Loads shift register with data[3:0] instead of shifting. // (3) ena: Shift right (q[3] becomes zero, q[0] is shifted out and disappears). // (4) q: The contents of the shift register. If both the load and ena inputs are asserted (1), the load input has higher priority. module top_module( input clk, input areset, input load, input ena, input [3:0] data, output reg [3:0] q);
always @(posedge clk or posedge areset) begin
if (areset) begin
// Asynchronous positive edge triggered areset
q <= 4’b0000;
end else if (load) begin
// Synchronous active high signals load
q <= data;
end else if (ena) begin
// Shift right
q <= {1’b0, q[3:1]};
end
end
endmodule
Listing 3: Correct Code
always @(posedge clk or posedge areset) begin
if (areset) begin
// Asynchronous positive edge triggered areset
q <= 4’b0;
end else if (load) begin
// Synchronous active high signals load
q <= data;
end else if (ena) begin
// Shift right
q <= {q[2:0], 1’b0};
end
end
endmodule
Listing 4: Error Code
LLM generated Error Report Error Type: shifting operation Category: Sequential: shift registers Description: The error in the erroneous implementation is related to the shifting operation in the shift register. q <= {q[2:0], 1’b0}; This line is intended to shift the bits in the register to the right, but it incorrectly connects the least significant bit (q[0]) to the most significant bit (q[3]) instead of connecting a zero to the most significant bit. This results in a circular shift instead of a right shift. To correct this error, the line should be changed to: q <= {1’b0, q[3:1]}; Here are the steps to repair the erroneous implementation 1. Identify the line of code that performs the shifting operation 2. Notice that the least significant bit (q[0]) is connected to the most significant bit (q[3]) instead of a zero. 3. Change the line of code to connect a zero to the most significant bit and shift the remaining bits to the right 4. Test the implementation to ensure that it behaves as expected.
LLM constructed Repair data from open-source code Input: You are provided with a Verilog module that implements a shift register. However, the implementation contains a common error related to shifting operations. Your task is to identify the error, which involves a mistake in the way bits are shifted within the register. Specifically, the problem results in an incorrect circular shift rather than the intended right shift operation. You need to fix the error to ensure that the register shifts correctly. Erroneous Implementation: module shift_registers_0 (clk, clken, SI, SO); parameter WIDTH = 32; input clk, clken, SI; output SO; reg [WIDTH-1:0] shreg; always @(posedge clk) begin if (clken) // Erroneous shifting operation: circular shift instead of right shift shreg = {shreg[WIDTH-2:0], shreg[WIDTH-1]}; end assign SO = shreg[WIDTH-1]; endmodule In this erroneous implementation: - The shift register incorrectly shifts the most significant bit (MSB) into the least significant bit (LSB), creating a circular shift. - The correct behavior should be shifting in the new input bit (‘SI’) while moving existing bits to the right. Hints for Fixing: 1. **Identify the shifting operation error:** The line ‘shreg = {shreg[WIDTH-2:0], shreg[WIDTH-1]};’ causes a circular shift because the MSB is incorrectly fed back into the LSB position. 2. **Fix the shifting logic:** Replace the erroneous MSB assignment with the correct new input bit (‘SI’). The correct line should be ‘shreg = {shreg[WIDTH-2:0], SI};’ to shift in the new input bit from the ‘SI‘ input and ensure the bits shift to the right. By fixing the shifting operation, the register will now shift correctly, with the new input bit filling the LSB and the existing bits shifting to the right. Output: module shift_registers_0 (clk, clken, SI, SO); parameter WIDTH = 32; input clk, clken, SI; output SO; reg [WIDTH-1:0] shreg; always @(posedge clk) begin if (clken) shreg = {shreg[WIDTH-2:0], SI}; end assign SO = shreg[WIDTH-1]; endmodule
Figure 8: Another example based on shift4 from VerilogEval-Human.
Problem description: multi_booth_8bit from RTLLM Please act as a professional verilog designer. Implement an 8-bit Radix-4 booth multiplier that performs the multiplication of two 8-bit inputs (a and b) using the Booth algorithm. It utilizes a clock signal (clk), and a reset signal (reset), and provides the product output (p) and a ready signal (rdy). The ready signal (rdy) is set to 1 to indicate the completion of the multiplication process. Module name: multi_booth_8bit Input ports: clk: Clock signal used for synchronous operation. reset: Reset signal used to initialize the multiplier module. a: 8-bit input representing the multiplicand. b: 8-bit input representing the multiplier. Output ports: p: 16-bit output representing the product of the multiplication. rdy: Ready signal indicating the completion of the multiplication operation. Implementation: On the positive edge of the clock signal (clk) or the positive edge of the reset signal (reset), the module performs the multiplication process. If the reset signal (reset) is high, two 16-bit registers multiplier <= 8a[7], a and multiplicand <= 8b[7], b. If the reset signal (reset) is low, indicating normal operation, the module checks if the counter (5bit ctr) is less than 16. If the counter (ctr) is less than 16, the multiplicand register (multiplicand) is left-shifted by 1 to simulate the Booth algorithm’s shifting operation. If the current bit of the multiplier register (multiplier[ctr]) is 1, indicating a positive Booth encoding, the product register (p) accumulates the value of the multiplicand register (multiplicand). The counter (ctr) is incremented by 1. Once the counter (ctr) reaches 16, indicating the completion of the multiplication process, the ready signal (rdy) is set to 1. Give me the complete code.
module multi_booth_8bit (p, rdy, clk, reset, a, b);
input clk, reset;
input [7:0] a, b;
output reg [15:0] p;
output reg rdy;
reg [15:0] multiplicand, multiplier;
reg [4:0] ctr;
always @(posedge clk or posedge reset) begin
if (reset) begin
multiplier <= {{8{a[7]}}, a};
multiplicand <= {{8{b[7]}}, b};
p <= 0;
ctr <= 0;
rdy <= 0;
end else begin
if (ctr < 16) begin
multiplicand <= multiplicand << 1;
if (multiplier[ctr]) begin
p <= p + multiplicand;
end
ctr <= ctr + 1;
end else begin
rdy <= 1;
end
end
end
endmodule
Listing 5: Correct Code
module multi_booth_8bit (p, rdy, clk, reset, a, b);
input clk, reset;
input [7:0] a, b;
output reg [15:0] p;
output reg rdy;
reg [15:0] multiplier, multiplicand;
reg [4:0] ctr;
always @(posedge clk or posedge reset) begin
if (reset) begin
multiplier <= {{8{a[7]}}, a};
multiplicand <= {{8{b[7]}}, b};
ctr <= 0;
rdy <= 0;
end else begin
if (ctr < 16) begin
multiplicand <= multiplicand << 1;
if (multiplier[ctr])
p <= p + multiplicand;
ctr <= ctr + 1;
end else begin
rdy <= 1;
end
end
end
endmodule
Listing 6: Error Code
LLM generated Error Report Error Type: Incorrect initialization Category: Sequential: counters Description: In the erroneous implementation, the p register is not initialized to zero during reset. This can lead to incorrect multiplication results, as the initial value of p is unknown and may contain random values. The correct implementation initializes the p register to zero during reset, ensuring that the multiplication result is accurate. To repair the erroneous implementation, follow these steps: 1. Locate the always block that handles the reset and clock edge. In this case, it’s the block that starts with always @(posedge clk or posedge reset). 2. Within the if (reset) block, add a line to initialize the p register to zero: p <= 0;.
LLM constructed Repair data from open-source code Input: Problem Description: You are given a Verilog module that implements a 4-bit up counter. The counter should reset to zero when the reset signal is active and increment with each positive clock edge. However, there is a common error in this implementation: the counter_up register is not correctly initialized during reset, which may lead to unpredictable behavior after reset. Your task is to identify and fix this initialization error. Erroneous Implementation: module up_counter( input clk, input reset, output [3:0] counter ); reg [3:0] counter_up; // up counter always @(posedge clk or posedge reset) begin if (reset) counter_up <= 4’d3; // Incorrect initialization value else counter_up <= counter_up + 4’d1; end assign counter = counter_up; endmodule In this erroneous implementation: * The counter_up register is incorrectly initialized to 4’d3 during reset, instead of being reset to zero (4’d0), which can cause incorrect counter behavior when reset is activated. Hints for Fixing: 1. Identify the initialization issue: The counter should reset to zero, but currently, it is being set to 4’d3 during reset. This causes incorrect counting behavior after a reset. 2. Fix the reset logic: Modify the if (reset) block in the always block to properly initialize the counter to zero. Change counter_up <= 4’d3; to counter_up <= 4’d0;. 3. Ensure proper reset behavior: After fixing the initialization, the counter will correctly reset to zero when the reset signal is active and increment correctly on each clock cycle thereafter. Output: module up_counter( input clk, input reset, output[3:0] counter ); reg [3:0] counter_up; // up counter always @(posedge clk or posedge reset) begin if(reset) counter_up <= 4’d0; else counter_up <= counter_up + 4’d1; end assign counter = counter_up; endmodule
Figure 9: Another example based on multi_booth_8bit from RTLLM.
Appendix C Examples of Correct-by-Construction Data for Non-Textual Representations
C.1 Karnaugh Maps and Truth Tables
Problem Implement the circuit described by the Karnaugh map below. // c // ab 0 1 // 00 | 1 | 0 // 01 | 0 | 0 // 11 | 0 | 0 // 10 | 0 | 0 module top_module( input a, input b, input c, output out ); Solution The input variables are: [’a’, b’, c’]. Based on the Karnaugh map, I can transform in to the following truth table: a | b | c | f 0 | 0 | 0 | 1 0 | 0 | 1 | 0 0 | 1 | 0 | 0 0 | 1 | 1 | 0 1 | 0 | 0 | 0 1 | 0 | 1 | 0 1 | 1 | 0 | 0 1 | 1 | 1 | 0 The minterms (when output is 1) are: (0,0,0) => (~a & ~b & ~c) This corresponds to the following minterms logic: ‘(~a & ~b & ~c)‘ Finally, based on the above logic equation, I can now write the Verilog code that could be described by the Karnaugh map: “‘ module top_module( input a, input b, input c, output out ); assign out = (~a & ~b & ~c); endmodule “‘
Figure 10: Example of a generated problem with Karnaugh Map representation in the problem description. We include detailed step-by-step analysis for reaching the correct solution.
C.2 State Transition Graphs and Tables
Problem Given the state-assigned table shown below, implement the logic functions Y[0] and z. // Present state y[2:0] | Next state Y[2:0] x=0, Next state Y[2:0] x=1 | Output z // 000 | 010, 011 | 1 // 001 | 100, 010 | 0 // 010 | 001, 100 | 1 // 011 | 011, 100 | 0 // 100 | 100, 001 | 0 Solution The state transition is as follows: // state | Next state in=0, Next state in=1 | Output // A | C, D | 1 // B | E, C | 0 // C | B, E | 1 // D | D, E | 0 // E | E, B | 0 The transition logic is then: A: next = x ? D : C; B: next = x ? C : E; C: next = x ? E : B; D: next = x ? E : D; E: next = x ? B : E; The output is 1 for states: A, C. Thus the output logic is: assign z = ( y == A || y == C ); Y0 corresponds to 001 (A), 011 (D). Finally, below is the Verilog code for the finite state machine: “‘ module top_module ( input clk, input x, input [2:0] y, output reg Y0, output reg z ); reg [2:0] next_state; parameter A=0, B=1, C=2, D=3, E=4; always_comb begin case(y) A: next_state = x ? D : C; B: next_state = x ? C : E; C: next_state = x ? E : B; D: next_state = x ? E : D; E: next_state = x ? B : E; default: next_state = x; endcase end assign z = ( y == A || y == C ); assign Y0 = ( next_state == B || next_state == D ); endmodule “‘
Figure 11: Example of problem with state transition tables.
Problem This is a Moore state machine with four states, four inputs, and one output. Implement this state machine in Verilog. Reset is an active-high synchronous reset to state D. // D (out=0) in0=1–> D // D (out=0) in0=0–> A // C (out=0) in1=1–> D // C (out=0) in1=0–> B // B (out=1) in2=1–> D // B (out=1) in2=0–> B // A (out=0) in3=1–> B // A (out=0) in3=0–> C Solution The finite state machine has four inputs, and the state transition logic is as follows: D: next = in0 ? D : A; C: next = in1 ? D : B; B: next = in2 ? D : B; A: next = in3 ? B : C; The output is 1 for states: B. Thus the output logic is: assign out = ( state == B );‘. Finally, below is the Verilog code for the finite state machine: “‘ module top_module ( input clk, input reset, input in0, input in1, input in2, input in3, output out ); parameter D=0, C=1, B=2, A=3; reg state; reg next; always_comb begin case(state) D: next = in0 ? D : A; C: next = in1 ? D : B; B: next = in2 ? D : B; A: next = in3 ? B : C; default: next = x; endcase end always @(posedge clk) begin if (reset) state <= D; else state <= next; end assign out = ( state == B ); endmodule “‘
Figure 12: Example of problem with Moore FSM.
Problem The following diagram is a Mealy machine. Implement in Verilog using one-hot encoding. Resets into state A and reset is asynchronous active-high. // A x=0 (z=0)–> D // A x=1 (z=1)–> C // B x=0 (z=1)–> C // B x=1 (z=0)–> B // C x=0 (z=0)–> C // C x=1 (z=0)–> D // D x=0 (z=1)–> C // D x=1 (z=0)–> B Solution From the transition diagram, we have the following transition logic: // state | next state in=0, next state in=1 // A | D, C // B | C, B // C | C, D // D | C, B Thus the state transition logic is as follows: A: next = x ? C : D; B: next = x ? B : C; C: next = x ? D : C; D: next = x ? B : C; The output is 1 for states: (A, x), (B, ~x), (D, ~x). Thus the output logic is: assign z = ( ( state == A & x ) || ( state == B & ~x ) || ( state == D & ~x ) );‘. Finally, below is the Verilog code for the finite state machine: “‘ module top_module ( input clk, input areset, input x, output z ); parameter A=2’b00, B=2’b01, C=2’b10, D=2’b11; reg [1:0] state; reg [1:0] next_state; always_comb begin case(state) A: next_state = x ? C : D; B: next_state = x ? B : C; C: next_state = x ? D : C; D: next_state = x ? B : C; default: next_state = x; endcase end always @(posedge clk, posedge areset) begin if (areset) state <= A; else state <= next_state; end assign z = ( ( state == A & x ) || ( state == B & ~x ) || ( state == D & ~x ) ); endmodule “‘
Figure 13: Example of problem with Mealy FSM.
Problem The following is the state transition table for a Moore state machine with one input, one output, and four states. Use the following one-hot state encoding: A=4’b0001, B=4’b0010, C=4’b0100, D=4’b1000. Derive state transition and output logic equations by inspection assuming a one-hot encoding. Implement only the state transition logic and output logic (the combinational logic portion) for this state machine. // state | Next state in=0, Next state in=1 | Output // A | B, A | 0 // B | B, C | 1 // C | D, A | 1 // D | B, C | 0 Solution Based on the state transition table, we can obtain the next state from observing the row (previous state) and column (input). Next state is A on the following (row, column): (A, in=1) (C, in=1). This correspond to the following logic: state[A] & in || state[C] & in‘. Next state is B on the following (row, column): (A, in=0) (B, in=0) (D, in=0). This correspond to the following logic: state[A] & ~in || state[B] & ~in || state[D] & ~in‘. Next state is C on the following (row, column): (B, in=1) (D, in=1). This correspond to the following logic: state[B] & in || state[D] & in‘. Next state is D on the following (row, column): (C, in=0). This correspond to the following logic: state[C] & ~in‘. The output is 1 for states: B, C. Thus the output logic is: assign out = ( state[B] || state[C] );‘. Finally, below is the Verilog code for the finite state machine: “‘ module top_module ( input in, input [3:0] state, output reg [3:0] next_state, output out ); parameter A=0, B=1, C=2, D=3; assign next_state[A] = state[A] & in || state[C] & in; assign next_state[B] = state[A] & ~in || state[B] & ~in || state[D] & ~in; assign next_state[C] = state[B] & in || state[D] & in; assign next_state[D] = state[C] & ~in; assign out = ( state[B] || state[C] ); endmodule “‘
Figure 14: Example of problem with in-edge focused solution shown in Figure 3.
C.3 Waveforms
Problem This is a combinational circuit. Read the simulation waveforms to determine what the circuit does, then implement it. // time a b c d q // 0ns 0 0 0 0 0 // 5ns 0 0 0 0 0 // 10ns 0 0 0 0 0 // 15ns 0 0 0 0 0 // 20ns 0 0 0 1 0 // 25ns 0 0 1 0 0 // 30ns 0 0 1 1 0 // 35ns 0 1 0 0 0 // 40ns 0 1 0 1 0 // 45ns 0 1 1 0 0 // 50ns 0 1 1 1 0 // 55ns 1 0 0 0 1 // 60ns 1 0 0 1 1 // 65ns 1 0 1 0 0 // 70ns 1 0 1 1 1 // 75ns 1 1 0 0 0 // 80ns 1 1 0 1 1 // 85ns 1 1 1 0 0 // 90ns 1 1 1 1 0 Solution Based on the simulation waveform, I can transform in to the following truth table: a | b | c | d | f 0 | 0 | 0 | 0 | 0 0 | 0 | 0 | 1 | 0 0 | 0 | 1 | 0 | 0 0 | 0 | 1 | 1 | 0 0 | 1 | 0 | 0 | 0 0 | 1 | 0 | 1 | 0 0 | 1 | 1 | 0 | 0 0 | 1 | 1 | 1 | 0 1 | 0 | 0 | 0 | 1 1 | 0 | 0 | 1 | 1 1 | 0 | 1 | 0 | 0 1 | 0 | 1 | 1 | 1 1 | 1 | 0 | 0 | 0 1 | 1 | 0 | 1 | 1 1 | 1 | 1 | 0 | 0 1 | 1 | 1 | 1 | 0 The minterms (when output is 1) are: (1,0,0,0) => (a & ~b & ~c & ~d) (1,0,0,1) => (a & ~b & ~c & d) (1,0,1,1) => (a & ~b & c & d) (1,1,0,1) => (a & b & ~c & d) This corresponds to the following minterms logic: ‘(a & ~b & ~c & ~d) | (a & ~b & ~c & d) | (a & ~b & c & d) | (a & b & ~c & d)‘ Finally, based on the above logic equation, I can now write the Verilog code: “‘ module top_module( input a, input b,
input c, input d, output q ); assign q = (a & ~b & ~c & ~d) | (a & ~b & ~c & d) | (a & ~b & c & d) | (a & b & ~c & d); endmodule “‘
Figure 15: Example of a waveform problem for combinational circuits.
Problem This is a sequential circuit. Read the simulation waveforms to determine what the circuit does, then implement it. // time clk reset in out // 0ns 0 1 0 x // 5ns 1 1 0 1 // 10ns 0 1 0 1 // 15ns 1 0 0 1 // 20ns 0 0 0 1 // 25ns 1 0 1 0 // 30ns 0 0 1 0 // 35ns 1 0 0 1 // 40ns 0 0 0 1 // 45ns 1 0 0 0 // 50ns 0 0 0 0 // 55ns 1 0 0 1 // 60ns 0 0 0 1 // 65ns 1 0 1 0 // 70ns 0 0 1 0 // 75ns 1 0 1 1 // 80ns 0 0 1 1 // 85ns 1 0 0 1 // 90ns 0 0 1 1 … Simulation waveform redacted … Solution From the waveform, we have the following transition logic and output logic: // state | Next state in=0, Next state in=1 | Output // D | C, B | 1 // C | B, D | 0 // B | D, A | 1 // A | C, B | 0 Thus the state transition logic is as follows: D: next = in ? B : C; C: next = in ? D : B; B: next = in ? A : D; A: next = in ? B : C; The output is 1 for states: D, B. Thus the output logic is: assign out = ( state == D || state == B );‘. Finally, below is the Verilog code for the finite state machine:
“‘ module top_module ( input clk, input in, input reset, output out ); parameter D=0, C=1, B=2, A=3; reg state; reg next; always_comb begin case(state) D: next = in ? B : C; C: next = in ? D : B; B: next = in ? A : D; A: next = in ? B : C; default: next = x; endcase end always @(posedge clk) begin if (reset) state <= D; else state <= next; end assign out = ( state == D || state == B ); endmodule “‘
Figure 16: Example of a waveform problem for sequential circuits.
Appendix D Prompt Templates
D.1 Synthetic Data Generation
D.1.1 Self-Instruct
Your goal is to create a high-quality Verilog problem. * Guidelines for designing the problem description: 1. This should be **completely self-contained**, providing all the contextual information one needs to understand and solve the problem. 2. Assume common verilog knowledge, but ensure that any specific context, variables, or code snippets pertinent to this problem are explicitly included. 3. Do not include the code snippet in the problem. 4. The problem should be desinged for the programmers to solve it with one verilog module. 5. The problem description section should be enclosed within <PROBLEM> </PROBLEM> tags. Now, Please use your creativity to create a brand new high-quality Verilog problem.
Figure 17: Prompt used to generate initial 50 seed problems for Self-Instruct.
Your goal is to create a high-quality Verilog problem. * Guidelines for designing the problem description: 1. This should be **completely self-contained**, providing all the contextual information one needs to understand and solve the problem. 2. Assume common verilog knowledge, but ensure that any specific context, variables, or code snippets pertinent to this problem are explicitly included. 3. Do not include the code snippet in the problem. 4. The problem should be desinged for the programmers to solve it with one verilog module. 5. The problem description section should be enclosed within <PROBLEM> </PROBLEM> tags. Below shows some examples: <PROBLEM> {seed problems} </PROBLEM> Now, Please use your creativity to create a brand new high-quality Verilog problem.
Figure 18: Prompt used for Self-Instruct.
D.1.2 OSS-Instruct
Your goal is to create a high-quality Verilog problem. * Guidelines for designing the problem description: 1. This should be **completely self-contained**, providing all the contextual information one needs to understand and solve the problem. 2. Assume common verilog knowledge, but ensure that any specific context, variables, or code snippets pertinent to this problem are explicitly included. 3. Do not include the code snippet in the problem. 4. The problem should be designed for the programmers to solve it with one Verilog module. * Guidelines for the problem description format: The problem description section should be enclosed within <PROBLEM> </PROBLEM> tags. Please increase the difficulty of the given programming test question a bit. You can increase the difficulty using, but not limited to, the following methods: 1. Your new problem should not be directly solved by the original code snippet. 2. You can also change the bit-width requiremnt, how to reset internal signals (if applicable), and whether the solution needs a clock signal (combinatorial versus sequential logic). If you do have a reset method that is synchronous to a clock, make sure to add the clock signal to the problem module input. 3. Add new constraints and requirements to the original problem, adding approximately 10 additional words. 4. Replace a commonly used requirement in the programming task with a less common and more specific one. 5. If the original problem can be solved with only a few logical steps, please add more reasoning steps. Now, Please gain inspiration from the following random code snippet to create a high-quality Verilog problem. Code snippet for inspiration: ‘‘‘ {code snippet} ‘‘‘ Output:
Figure 19: Prompt used for OSS-Instruct. We also include prompts inspired from Evol-Instruct (Luo et al., 2024b) to increase problem difficulty.
D.1.3 Docu-Instruct
Your goal is to create a high-quality Verilog problem. * Guidelines for designing the problem description: 1. This should be **completely self-contained**, providing all the contextual information one needs to understand and solve the problem. 2. Assume common verilog knowledge, but ensure that any specific context, variables, or code snippets pertinent to this problem are explicitly included. 3. Do not include the code snippet in the problem. 4. The problem should be designed for the programmers to solve it with one Verilog module. * Guidelines for the problem description format: The problem description section should be enclosed within <PROBLEM> </PROBLEM> tags. Now, Please gain inspiration from the following textbook or wikipedia snippet to create a high-quality Verilog problem. The information might not be directly related to Verilog, but try to be make the problem as relevant as possible to the textbook issue discussed. Textbook snippet for inspiration: ‘‘‘ {document snippet} ‘‘‘ Output:
Figure 20: Prompt used for Docu-Instruct with Wikipedia and textbooks.
I am going to give you a concept and some descriptions about that concept. Based on the descriptions and concept name, determine if the concept belongs to one of the following categories: - Hardware description and modeling in Verilog. - Fundamental constructs such as modules, ports, and wires specific to Verilog. - Synthesis and optimization techniques employed in hardware design using Verilog. - Simulation tools and methodologies for verifying Verilog-based hardware designs. - Common design patterns and best practices in Verilog for efficient hardware implementation. - Programming concepts like loops, functions related to Verilog. - Hardware related concepts such as finite state machines that could be implemented in Verilog. - Algorithms that could be implemented in hardware, such as Fourier Transforms. Concept: {Wikipedia title} Description: {Wikipedia content} Do not make assumptions and only respond “Yes” if you are certain that the {Wikipedia title} is related to hardware design or Verilog coding language. Your answer should start with “Yes” or “No”.
Figure 21: Prompt used to filter Verilog related Wikipedia pages.
D.1.4 Non-textual Representations
Your goal is to create a high-quality Verilog problem. Specifically, we would like to test the skills of understanding Karnaugh maps and state transition diagrams. The problem description section should be enclosed within <PROBLEM> </PROBLEM> tags. Now, please gain inspiration from the following random code snippet to create a high-quality Verilog problem. Remember that the problem you generated must include Karnaugh maps in the format above. The random code snippet MUST be related to the solution. Your problem statement should be short and succinct (no more than 5 sentences) and you MUST generate a Karnaugh map in the problem description. Your problem description should not describe the Karnaugh map in words and should assume that the student need to decipher the Karnaugh map to solve the problem. Code snippet for inspiration: ‘‘‘ {code snippet} ‘‘‘ Below are two examples on how to represent Karnaugh map related questions in purely textual format. You should NOT use the following to generate the problem but only consider the style. <PROBLEM> Given the state-assigned table shown below, implement the finite-state machine. Reset should synchronous active high reset the FSM to state 000. // Present state y[2:0] | Next state y[2:0] x=0, Next state y[2:0] x=1, Output z // 000 | 000, 001 | 0 // 001 | 001, 100 | 0 // 010 | 010, 001 | 0 // 011 | 001, 010 | 1 // 100 | 011, 100 | 1 </PROBLEM> <PROBLEM> Implement the circuit described by the Karnaugh map below. // a // bc 0 1 // 00 | 0 | 1 | // 01 | 1 | 1 | // 11 | 1 | 1 | // 10 | 1 | 1 | </PROBLEM>
Figure 22: An prompt example to encourage LLMs to generate questions with Karnaugh Maps.
D.1.5 Prompts for Sampling Solutions with LLM Generated Problems